Find the pH and Concentration of H2SO3, HSO3-, and (SO3)2-

a) 0.050M H2SO3

b) 0.050M NaHSO3

c) 0.050M Na2SO3
Answer #1

(a). ICE table of H2SO3 is :

.................H2SO3............+...............H2O ---------------------> HSO3-.................+................H3O+

Initial (I).....0.050 M.................................................................0.0 M...................................0.0 M

Change(C)...- y ......................................................................+ y.........................................+ y

Equilibrium (E) ..(0.050-y) M....................................................y M.........................................y M

Expression of Ka is :

Ka = [HSO3-].[H3O+] / [H2SO3]

1.3 x 10-2 = y2/(0.050-y)

y2 + 0.013 y - 0.00065 = 0

On solving, we have

y = 0.0198

So,

[H3O+] = y = 0.0198 M

pH = - log [H3O+]

pH = - log 0.0198

pH = 1.70

---------------------------------------------------------------

(b).

ICE table of HSO3- is :

.................HSO3-............+...............H2O ---------------------> SO32-.................+................H3O+

Initial (I).....0.050 M.................................................................0.0 M...................................0.0 M

Change(C)...- y ......................................................................+ y.........................................+ y

Equilibrium (E) ..(0.050-y) M....................................................y M.........................................y M

Expression of Ka2 is :

Ka2 = y2/0.050-y

y<<<0.050, so neglect y as compare to 0.050, we have

6.3 x 10-8 x 0.050 = y2

y2 = 0.315 x 10-8

y = 5.61 x 10-5 = [ H3O+ ]

pH = - log 5.61 x 10-5

pH = 4.25

---------------------------------------------------------------------------------------------

(c). Kb1 of SO32- = 1.0 x 10-14/6.3 x 10-8 = 1.587 x 10-7

ICE table of SO32- is :

.................SO32-............+...............H2O ---------------------> HSO3-.................+................OH-

Initial (I).....0.050 M.................................................................0.0 M...................................0.0 M

Change(C)...- y ......................................................................+ y.........................................+ y

Equilibrium (E) ..(0.050-y) M....................................................y M.........................................y M

Expression of Kb1 is :

Kb1 = y2/0.050-y

y<<<0.050, so neglect y as compare to 0.050, we have

1.587 x 10-7 x 0.050 = y2

y2 = 7.93 x 10-9

y = 8.90 x 10-5 = [OH-]

pOH = - log[OH-]

pOH = -log 8.90 x 10-5

pOH = 4.05

pH = 14-pOH

pH = 14-4.05

pH = 9.95